OpenStudy (anonymous):

Write an equivelent expression: Use the Associative property of multipication to rewrite -5x(2r) another way

OpenStudy (anonymous):

The Associative Property of Multiplication is basically;\[4\times\left( 2\times3 \right)=\left( 4\times2 \right)\times3\] So in your problem it would be; \[-5x(2r)=(-5x \times2)r\]